Multiply the binomials. (2.5l – 0.5m) and (2.5l + 0.5m)
Please solve andanswer fast i need the answer for holiday homework.

Grade:8

2 Answers

Ayush 8A 7239
2 years ago
(2.5l – 0.5m)(2.5l+ 0.5m)
By using identity (a-b)(a+b) = a2 - b2
 
(2.5l – 0.5m)(2.5l+ 0.5m) = (2.5l)2 – (0.5m)2
= 6.25 l- 0.25m2
Ayush 8A 7239
2 years ago
By the way i have answeredall your questions which you asked i don’t why but your questions were comig to be answered in my questions page.
That’s why i solved them